OpenStudy (anonymous):

You are planning to build a mini hydro power plant. A river near your home is w = 4.0 m wide and h = 1.0 m deep. You measure the water flow to be v = 1.50 m/s over the brink of a waterfall H = 10.0 m high. The company that makes the turbine advertises e = 25% efficiency in converting the potential energy of the water into electric energy. Derive an expression for the generated power in terms of the above variables and any physics constants. How many 100watts bulbs will this power handle? You are planning to build a mini hydro power plant. A river near your home is w = 4.0 m wide and h = 1.0 m deep. You measure the water flow to be v = 1.50 m/s over the brink of a waterfall H = 10.0 m high. The company that makes the turbine advertises e = 25% efficiency in converting the potential energy of the water into electric energy. Derive an expression for the generated power in terms of the above variables and any physics constants. How many 100watts bulbs will this power handle? @Physics

OpenStudy (anonymous):

Can you please help me?

OpenStudy (anonymous):

4 x 1 x 1.5 = 6 cubic meters of water per second = 6000 liters of water = 6000 kilos of water amount of potential energy = m*g*h E = 6000 * 9.81 * 10 = 588600 joules 25% efficiency => 147150 joules of energy 1 Watt = 1 joule / second 147150 / 100 = 1471.5 bulbs

OpenStudy (anonymous):

ok thanks!:)

OpenStudy (anonymous):

can you pls help me with my next question as well?

OpenStudy (anonymous):

who knows ;)

OpenStudy (anonymous):

make a new thread tho

OpenStudy (anonymous):

idid

OpenStudy (anonymous):

its the one that starts off with Mike Harris...